In a survey of men in a certain country? (ages 20minus??29), the mean height was 63.5 inches with a standard deviation of 2.6 inches. ?(a) What height represents the 99th ?percentile? ?(b) What height represents the first? quartile?
Answer #1

(a). It is given that

\mu = 63.5

\sigma = 2.6

The z-score with a left tail of 99% is 2.33

z= \frac{X-\mu}{\sigma}

2.33= \frac{X-63.5}{2.6}

X= (2.33)(2.6) + 63.5 = 69.558

(b). For first quartile, that is, z-score for a left tail of 25%,

z = -0.6745

z = \frac{X-\mu}{\sigma}

-0.6745 = \frac{X - 63.5}{2.6}

X = 63.5 - (0.6745)(2.6) = 63.5 - 1.7537 = 61.7463 inches